# From Learning to Safety: A Direct Data-Driven Framework for Constrained Control

### v1

@@ 1.1 Background
In this paper, we focus on designing a direct data-driven safety filter (3DSF) based on our previously proposed concept of state-action control barrier functions (SACBFs) [23]. Unlike classical safety certificates, which only evaluate safety in the state space, the SACBF framework enables safety evaluation for each state-action pair. This key feature eliminates the need for explicit system dynamics during policy modification, making it particularly suitable for model-free learning-based control. The comparison between the proposed 3DSF and the existing indirect data-driven counterparts is illustrated in Fig. 1. For the indirect data-driven methods, both system identification and certificate learning should be performed one by one or simultaneously. The identified model and learned certificate are both used in the safety filter design. In comparison, the proposed 3DSF only uses an SACBF. A qualitative comparison of our method with others can be found in Section VII.
@@ 1.2 Contributions
(1) Optimization-based direct data-driven safe control: We propose a novel safe control framework for general nonlinear systems with nonlinear constraints. The main advantage is that the safe controller can be trained and implemented using state transition data only, without the need for system identification. The control framework incorporates a novel safety filter, which we call the direct data-driven safety filter (3DSF). This framework, which can be integrated with arbitrary learning-based controller synthesis methods, also provides formal guarantees on constraint satisfaction.
(2) Learning-based synthesis of SACBFs: We develop three distinct learning-based methods for synthesizing SACBFs directly from data: including SL from a known CBF, learning from an expert safe controller, and RL (self-learning).
(3) Robustness analysis via error-to-state safety (ESSf): We propose a systematic framework, called Error-to-State Safety (ESSf), to analyze how learning-induced errors in SACBFs affect overall safety performance. The framework motivates the approach of state constraint tightening followed by SACBF constraint relaxation to ensure that the learned SACBF-regulated controller meets safety requirements.
(4) A unified framework extending RL to constrained control: Building on the proposed optimization-based safe control, we present a general strategy to extend classical unconstrained value-based RL algorithms to constrained ones so that suboptimal performance regarding other tasks is achieved. This reversely shows another benefit of our proposed optimization-based approach: it separates the learning of the controller into two components: optimizing performance through learning the objective function, and ensuring safety through learning the constraints.
Additionally, for contribution (2), if system nonlinearities are known, we formulate the synthesis of a quadratic SACBF as solving a convex optimization problem with linear matrix inequality (LMI) constraints.
@@ 2.1 Notations
The set $\mathbb{R}$ ($\mathbb{R}_{\geq 0}$) is the set of (nonnegative) real numbers, and the set of real vectors with the dimension $n$ is denoted by $\mathbb{R}^{n}$. The sets $\mathbb{N}$ and $\mathbb{N}^{+}$ represent the set of non-negative integers and the set of positive integers, respectively. Besides, $\mathbb{N}_{a}=\{0,1,...,a\}$, and $\mathbb{N}^{+}_{a}=\{1,...,a\}$ for any positive integer $a$. The matrix $I_{n}$ is the identity matrix with the dimension $n\times n$. The notation $A^{\dagger}$ represents the right inverse of the matrix $A$. The relation $A\succeq 0$ means that the matrix $A$ is positive semi-definite. The determinant of a matrix $A$ is denoted by $\mathrm{det}(A)$. The dimension of a vector $x$ is denoted by $\mathrm{dim}(x)$. The norm $||x||_{2}$ is the Euclidian norm of the vector $x$. The number of elements in a finite set $S$ is called its cardinality, denoted by $|S|$. The set $\mathcal{B}_{\epsilon}(\bar{z}):=\{z\in\mathbb{R}^{n_{z}}\mid||z-\bar{z}||_{2}\leq\epsilon\}$ denotes the closed ball around $\bar{z}$ with radius $\epsilon$. The unit step function $\mathrm{step}(\cdot)$ returns 1 if the input is larger than 0 and returns 0 otherwise. A continuous function $\alpha(\cdot):[0,a)\to[0,\infty)$ for some $a\in(0,\infty]$ is said to belong to class $\mathcal{K}$ if it is strictly increasing and $\alpha(0)=0$. A class $\mathcal{K}$ function $\alpha$ is said to belong to class $\mathcal{K}_{\infty}$ if it further satisfies $a=\infty$ and $\alpha(r)\to\infty$ as $r\to\infty$.
2.2 Preliminaries
@@
Given an initial state $x_{0}$, we are interested in designing a deterministic control policy $\pi:\mathcal{X}\to\mathcal{U}$ to steer the trajectory of the system to a non-empty target region $X_{\mathrm{tar}}\subseteq X$. To achieve the control objective, for any state $x$ and input $u$, a stage cost is defined by $g\left(x,u\right)$, which is non-negative and upper-bounded. For any policy $\pi$, the value function $J^{\pi}\left(\cdot\right):\mathcal{X}\to[0,\infty)$ is defined by
@@
where $\gamma\in(0,1)$ is a discount factor. The objective is to find an optimal deterministic policy $\pi^{*}(\cdot):\mathcal{X}\to\mathcal{U}$ that solves the following infinite-horizon optimal control problem:
@@ 2.3 Control barrier functions for safety
Definition 1 (Control barrier function [26]).
@@
where $Q:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is a problem-dependent objective function. For instance, one can specify $Q$ as $||u-\pi^{\mathrm{unsafe}}(x)||_{2}$ where $\pi^{\mathrm{unsafe}}$ is a reference policy that may have some other acceptable performance. With this specification, (2.3) works as a safety filter to refine the potentially unsafe policy $\pi^{\mathrm{unsafe}}$. One can also let $Q$ be an approximation of the state-action optimal value function (Q function) of (3), that is,
@@
A valid CBF is sufficient to guarantee the safety of $\pi$ [4]. However, the main limitation of (2.3) is that the knowledge of $f$ is required. Even though some data-driven methods exist for learning CBFs from black-box models [17], system identification is necessary to implement (2.3). This limitation significantly restricts the application of safe filters and CBFs in reinforcement learning algorithms that do not require an explicit model. Moreover, the mismatch between the identified model and the real system, along with inaccuracies in learning CBFs, jointly affect the safety performance of $\pi$.
3 State-action CBFs and control parameterization
$\pi(x):=\arg\min_{u\in U}$ $\;Q(x,u)$
$\mathrm{s.t.}\;$ $Q^{B}(x,u)\leq 0,$ (7)
where $Q$ can be designed similarly to that in (2.3), and another function $Q^{B}(\cdot,\cdot):\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is used to add a constraint to regulate the behavior of $\pi$. To make $\pi$ satisfy the safety requirements $h(x_{t})\leq 0,\;\forall t\in\mathbb{N}$, we consider the definition of state-action CBFs, which we have introduced in [23].
@@ 3 State-action CBFs and control parameterization
There are two main features that distinguish SACBFs from standard CBFs. First, the SACBF is a function of both the states and the input. It not only characterizes the energy of a state but also quantifies the quality (safety) of selecting an input in a given state. Besides, the explicit form of the safe set is not specified in Definition 2.
According to the definition of SACBFs, we have the following properties for (3):
Lemma 1 (Safety of SACBFs[23]).
@@ 4 Learning state-action control barrier functions
In this section, we propose three learning-based approaches to approximate an SACBF $Q^{B}_{\omega}$. For the first one, we assume the knowledge of a valid CBF and use SL to obtain an SACBF. The second approach, inspired by [11], employs sampling-based methods to approximate the solution of a robust optimization problem. Such a solution is guaranteed to lead to a valid SACBF. This approach requires prior knowledge of a safety control policy (usually conservative w.r.t. task performance). For the third approach, we connect the synthesis of SACBFs with Hamilton-Jacobi reachability [27], and thereby propose an RL-based approach to obtain $Q^{B}_{\omega}$. The third approach only relies on the availability of state transition data. Besides, when the nonlinearity of the system is known, we also find that the synthesis of SACBFs can be achieved by solving a convex optimization problem with LMI conditions.
4.1 Supervised learning SACBFs from CBFs
We first present our simplest learning-based method for obtaining an SACBF, based on a rather restrictive assumption that we have a CBF at hand. In certain applications such as adaptive cruise control [4], CBFs can be manually designed based on state constraints, such as the distance between adjacent vehicles. However, as noted in the introduction, such CBFs cannot be used to design a safety filter (2.3) when the model is not fully known. In this subsection, we explore connections between SACBFs and CBFs and demonstrate that one can easily learn an SACBF from a given CBF.
Assumption 1.
A CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is known.
@@ 4.2 Learning SACBFs from an expert controller
In the presence of both state and input constraints, manually crafted CBFs often fail. Towards the goal of learning a valid SACBF, like in [11] where CBFs are synthesized from expert demonstrations, we assume the availability of an expert safe controller $\pi_{\mathrm{s}}(\cdot):\mathcal{X}\to\mathcal{U}$ in this subsection. This assumption is reasonable in certain scenarios, such as human-controlled systems (e.g., driving), where safety is achieved implicitly without an explicit model.
@@
There exists a continuous policy $\pi_{\mathrm{s}}$ and a compact set $\mathcal{S}_{0}\subseteq X$ ($\mathcal{S}_{0}$ can be unknown) such that with the initial condition $x_{0}\in\mathcal{S}_{0}$, the state-input trajectories of the system (1) with $\pi_{\mathrm{s}}$ always stay in $\mathcal{S}_{0}\times U$ and such that the states reach $X_{\mathrm{tar}}$ in a finite number of time steps $T$.
@@
Under Assumption 2, there exists a $\beta\in[0,1)$ such that the problem (9) is feasible and such that any optimal solution pair $(Q^{B*},q^{*})$ ensures that $Q^{B*}$ is an SACBF with the safe set $\mathcal{S}^{*}_{Q}=\{x\in\mathcal{X}|q^{*}(x)\leq 0\}$.
@@
$\mathcal{S}_{0}$ =$=\left\{x\in\mathcal{X}|q(x)\leq 0\right\},$
$\partial\mathcal{S}_{0}$ =$=\left\{x\in\mathcal{X}|q(x)=0\right\},$
$\operatorname{Int}(\mathcal{S}_{0})$ =$=\left\{x\in\mathcal{X}|q(x)<0\right\}$ (10)
@@
With the above definitions, we follow the proof of the Converse CBF Theorem in the continuous time domain [4, Proposition 3] to prove the feasibility of (9). In particular, for any continuous $q$ such that (9b) and (4.2) hold, define a function $\alpha(\cdot):[0,\infty)\to\mathbb{R}$ by
@@
When $x$ is such that $q(x)=0$, i.e., $x\in\partial\mathcal{S}_{0}$, by the invariance of $\mathcal{S}_{0}$, we have $q(f(x,\pi_{\mathrm{s}}(x)))\leq 0$, and consequently $\alpha(0)\leq 0$. Meanwhile, by taking the extreme values 00 and $-r$ of $q(f(x,\pi_{\mathrm{s}}(x)))$ and $q(x)$ respectively, we have $\min_{r>0}\frac{\alpha(r)}{r}\leq\min_{r>0}\frac{0-(-r)}{r}=1$, which, combined with $\alpha(0)\leq 0$, implies that there exists a class $\mathcal{K}$ function $\bar{\alpha}$ upper-bounding $\alpha$ and satisfying
@@
[…]
@@
For the optimization problem (9), we replace the infinite sets $\mathcal{S}_{0}$, $\mathcal{X}$, and $\mathcal{S}_{0}\times U$ by the finite sets $\mathcal{S}_{\mathrm{s}}$, $\mathcal{X}_{\mathrm{s}}$, and $\mathcal{S}_{\mathrm{s}}\times U_{\mathrm{s}}$, respectively. Here $U_{\mathrm{s}}$ is the set of inputs sampled in $U$ through some sampling strategy. As a result, (9) is simplified to a constrained optimization problem with a finite number of constraints. If $Q^{B}_{\omega}$ and $q$ are linear in $\omega$, the problem (9) becomes linear and can be solved efficiently. In a more general case such as when these functions are represented by deep neural networks, a tractable solution involves transforming the constraints into soft penalties and adding them into the objective.
@@
Due to the above constraint relaxation, the validity of the approximated solution is however not guaranteed. To get a valid SACBF, we can follow the approach of [11] to tighten the constraints in (9). Assuming $Q^{B}_{\omega}$, $q_{\omega}$, $\pi_{\mathrm{s}}$, and $f$ to be Lipschitz as well as having sufficient sampling, $Q^{B}_{\omega}$ will be a valid SACBF222“Sufficient sampling” means that $\mathcal{X}_{\mathrm{s}}\times U_{\mathrm{s}}$ constitutes an $\epsilon$-net of $\mathcal{X}\times U$ with $\epsilon>0$ [11]. Namely, $\forall x\in\mathcal{X}$ and $\forall u\in U$, $\exists x^{\prime}\in\mathcal{X}_{\mathrm{s}}$ and $\exists u^{\prime}\in U_{\mathrm{s}}$ such that $||x-x^{\prime}||^{2}_{2}+||u-u^{\prime}||^{2}_{2}\leq\epsilon^{2}$.. Although this constraint tightening approach will guarantee that $Q^{B}_{\omega^{*}}$ is a valid SACBF, it may be conservative under insufficient sampling. Furthermore, estimating the Lipschitz constants is a non-trivial task. To address this limitation, we will propose a new tightening approach exploits the inherent robustness of the SACBF. This will be elaborated on in Section 5.
4.3 Learning SACBFs via RL
In the previous subsection, we have proposed an optimization-based approach to learn SACBFs from a known safe controller. This approach has two main limitations. First, a safe controller is sometimes hard to obtain as a prior. Second, the safe set of the learned SACBF cannot be larger than the safe region $\mathcal{S}_{0}$ of the safe controller. On the other hand, it is desirable to learn an SACBF with the largest possible safe set directly from randomly generated transition data, without any expert supervision. For this purpose, we observe that the proposed SACBFs show some similarities with state-action value functions (Q functions) in RL. These similarities motivate us to use RL methods to synthesize SACBFs.
We consider the optimal value function of the following optimal control problem:
$B^{*}(x):=$ $\min_{\{(x_{t},u_{t})\}_{t=0}^{\infty}}\max_{t\in\mathbb{N}}h(x_{t})$
$\quad\mathrm{s.t.}\;(1),\;\;u_{t}\in U,\;t\in\mathbb{N},\text{ and }x_{0}=x,$ (15)
which is a typical reachability problem [27] in the discrete-time domain.

### v2

@@ 1.1 Background
In this paper, we focus on designing a direct data-driven safety filter (3DSF) based on our previously proposed concept of state-action control barrier functions (SACBFs) [24]. Unlike classical safety certificates, which only evaluate safety in the state space, the SACBF framework enables safety evaluation for each state-action pair, and the safety condition is enforced at the current state and action without propagating the dynamics forward. As a result, the corresponding safety filter can assess the safety of candidate control actions without explicitly querying or simulating the system dynamics. The comparison between the proposed 3DSF and the existing indirect data-driven counterparts is illustrated in Fig. 1.
@@ 1.2 Contributions
(1) We propose a direct data-driven safe control framework based on SACBFs, which combines the advantages of learning-based control (adaptivity to model-free cases) and optimization-based control (explicit inclusion of safety constraints). The core mechanism relies on our previously introduced safety certificate called the SACBF [24], for which we develop three distinct data-driven synthesis methods, including SL from a known CBF, learning from an expert safe controller, and RL. The framework can be trained and implemented using only state transition data, without system identification, and can integrate with various learning-based controller synthesis methods while providing formal safety guarantees.
(2) To analyze how learning-induced errors in SACBFs affect overall safety performance, we propose a robust analysis method, called Error-to-State Safety (ESSf). The analysis motivates the approach of state constraint tightening followed by SACBF constraint relaxation to ensure that the learned SACBF-based controller meets safety requirements.
(3) To enable near-optimal control performance, building on the proposed framework, we present a general strategy to extend classical unconstrained value-based RL algorithms to constrained ones. This design highlights an important feature of our approach: it separates the learning of the controller into two components: optimizing performance through learning the objective function, and ensuring safety through learning the constraints.
@@ 2.1 Notations
[…]
2.2 Problem formulation
@@
For any initial state $x\in\mathcal{X}$ and any deterministic control policy $\pi:\mathcal{X}\to\mathcal{U}$, we define an infinite-horizon cost $J^{\pi}\left(\cdot\right):\mathcal{X}\to[0,\infty)$:
@@
where the stage cost $g$ is non-negative and upper-bounded and $\gamma\in(0,1)$ is a discount factor. The objective is to find an optimal deterministic policy $\pi^{*}(\cdot):\mathcal{X}\to\mathcal{U}$ that solves the following infinite-horizon optimal control problem:
@@ 2.3 Control barrier functions for safety
Definition 1 (Discrete-time control barrier function [1]).
@@
where $Q:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is a problem-dependent objective function. For instance, one can specify $Q$ as $u-\pi^{\mathrm{unsafe}}(x)||_{2}$ where $\pi^{\mathrm{unsafe}}$ is a reference policy that may have some other acceptable performance. With this specification, (2.3) works as a safety filter to refine the potentially unsafe policy $\pi^{\mathrm{unsafe}}$. One can also let $Q$ be an approximation of the state-action optimal value function (Q function) of (3), that is,
@@
A valid CBF is sufficient to guarantee the safety of $\pi$ [2]. However, the main limitation of CBFs is that the knowledge of $f$ is required. Even though many data-driven methods exist for learning CBFs from black-box models, such as learning CBFs by temporal-difference learning [46] and by policy evaluation [45, 55], system identification is necessary to implement the CBF-based safety filter (2.3). This limitation significantly restricts the application of existing safe filters and CBFs in RL algorithms that do not require an explicit model. Moreover, the mismatch between the identified model and the real system, along with inaccuracies in learning CBFs, jointly affect the safety performance of $\pi$.
2.4 State-action control barrier function
wWe consider the definition of state-action CBFs, which we have introduced in [24].
@@ 2.4 State-action control barrier function
Unlike a standard CBF, which evaluates the safety level of a state, the SACBF is a function of both the states and the input. This additional dependence on the input allows the SACBF to not only assess whether a state is safe, but also to quantify the safety margin associated with taking a particular action at that state.
3 Control parameterization
$\pi(x):=\arg\min_{u\in U}$ $\;Q(x,u)$
$\mathrm{s.t.}\;$ $Q^{B}(x,u)\leq 0,$ (7)
where $Q$ can be designed similarly to that in (2.3), and $Q^{B}(\cdot,\cdot):\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is an SACBF.
We have the following properties for (3):
Lemma 1 (Safety of SACBFs).
@@ 3 Control parameterization
The proof can be obtained by a straightforward adaptation of the proof of [24, Lemma 1].
@@ 4 Learning state-action control barrier functions
In this section, we present three approaches that differ in their assumptions and the strength of the resulting safety guarantees. We begin with the most general method (RL approach), which does not require any assumptions. This method relies on off-the-shelf value-based RL to approximate $Q^{B}$. However, the learning error may in principle become unbounded. Consequently, the RL approach cannot offer formal safety guarantees. Motivated by this limitation, we then introduce two additional approaches that impose stronger assumptions. The second approach, inspired by [44], employs sampling-based methods to approximate the solution of a robust optimization problem. Such a solution is guaranteed to lead to a valid SACBF. This approach requires prior knowledge of a safety control policy. For the last one, we assume the knowledge of a valid CBF and use SL to obtain an SACBF. These assumptions restrict generality but allow the approximation errors to be explicitly characterized, enabling us to derive provable and rigorous safety guarantees.
4.1 Learning SACBFs via RL
We observe that the proposed SACBFs show some similarities with state-action value functions (Q functions) in RL. These similarities motivate us to use RL methods to synthesize SACBFs. We consider the optimal value function of the following optimal control problem:
$B^{*}(x):=$ $\min_{\{(x_{t},u_{t})\}_{t=0}^{\infty}}\max_{t\in\mathbb{N}}h(x_{t})$
$\quad\mathrm{s.t.}\;(1),\;\;u_{t}\in U,\;t\in\mathbb{N},\text{ and }x_{0}=x,$ (8)
which is a typical reachability problem [18] in the discrete-time domain. The optimal value function $B^{*}$, which can be obtained by dynamic programming, is a CBF with the safe set being the maximal control-invariant set [12].
@@ 4.1 Learning SACBFs via RL
Using Proposition 1 and letting $\beta_{B}=1$, we know that the state-action optimal value function of (4.1), defined by
$Q^{B*}(x,u):=\max\{h(x),\;B^{*}(f(x,u))\},$ (10)
$Q^{B}(x,u)=\max\{h(x),\min_{u^{+}\in U}Q^{B}(f(x,u),u^{+})\}.$ (11)
Proposition 2.
Consider the Bellman optimality equation (11). Any solution $Q^{B}$ satisfying $\min_{x\in\mathcal{X},u\in U}Q^{B}(x,u)\leq 0$ is an SACBF with the safe set $\mathcal{S}_{Q}=\{x\in\mathbb{R}^{n_{x}}|\min_{u\in U}Q^{B}(x,u)\leq 0\}$. Moreover, $Q^{B*}$ is the smallest solution, i,e., $Q^{B*}(x,u)\leq Q^{B}(x,u),\;\forall(x,u)\in\mathcal{X}\times U$ holds for any solution $Q^{B}$.
Proof.
It follows from $\min_{x\in\mathcal{X},u\in U}Q^{B}(x,u)\leq 0$ that $\mathcal{S}_{Q}$ is non-empty. From (11), we deduce that $h(x)\leq\min_{u\in U}Q^{B}(x,u)\leq 0,\;\forall x\in\mathcal{S}_{Q}$. The above arguments prove that $Q^{B}$ satisfies (i) of Definition 2. The second condition of Definition 2 directly follows from $\mathcal{S}_{Q}=\{x\in\mathbb{R}^{n_{x}}|\min_{u\in U}Q^{B}(x,u)\leq 0\}$. Finally, if $Q(x,u)\leq 0$, by (11) we have $\min_{u^{+}\in U}Q^{B}(f(x,u),u^{+})\leq 0$, which means that the successor state $f(x,u)\in\mathcal{S}_{Q}$.
To prove that $Q^{B*}$ is the smallest solution, consider any solution $Q^{B}$. It follows from (11) that
$h(x)\leq Q^{B}(x,u),\;\forall(x,u)\in\mathcal{X}\times U.$ (12)
Applying the Bellman operation $\Gamma_{B}(\cdot):=\max\{h(x),\min_{u^{+}\in U}(\cdot)(f(x,u),u^{+})\}$ to both sides of (12) infinitely many times and noting that $\Gamma_{B}Q^{B}=Q^{B}$ and $\Gamma^{\infty}_{B}h=Q^{B*}$, we obtain that $Q^{B*}(x,u)\leq Q^{B}(x,u),\;\forall(x,u)\in\mathcal{X}\times U$. This completes the proof of Proposition 2. ∎
Remark 2.
Since the reachability problem is undiscounted, most RL algorithms cannot, in general, guarantee convergence to $Q^{B*}$. Tractable approaches to alleviate this issue include introducing a discount factor [18, 45] and considering the reachability to a known invariant set [11]. We do not employ these approaches because they either flatten the value function in the safe set or require prior knowledge of an invariant set, which is restrictive in the model-free setting. Addressing this challenge will be a topic for future work.
@@ 4.2 Learning SACBFs from an expert controller
The RL approach usually does not have performance guarantees in theory. Towards the goal of learning a valid SACBF, like in [44] where CBFs are synthesized from expert demonstrations, we assume the availability of an expert safe controller $\pi_{\mathrm{s}}(\cdot):\mathcal{X}\to\mathcal{U}$ in this subsection.
@@
There exists a continuous policy $\pi_{\mathrm{s}}$ and a compact set $\mathcal{S}_{0}\subseteq X$ ($\mathcal{S}_{0}$ can be unknown) such that with the initial condition $x_{0}\in\mathcal{S}_{0}$, the state-input trajectories of the system (1) with $\pi_{\mathrm{s}}$ always stay in $\mathcal{S}_{0}\times U$ and such that the states reach a non-empty target set $X_{\mathrm{tar}}\subseteq X$ in a finite number of time steps $T$.
Note that we do not require the explicit form of the safe policy. This assumption is reasonable in certain scenarios, such as human-controlled systems (e.g., driving), where safety is achieved by feedback on state variables (e.g., position and velocity) without an explicit model. Other types of expert safe controllers that do not require model knowledge include artificial potential fields (APF) [51] for obstacle avoidance and safe controllers derived from Hamilton-Jacobi reachability, which have recently been extended to model-free settings [18]. The target set is introduced to determine whether the sampled trajectories are safe (i.e., satisfy the constraints and reach the target) or not. Once the trajectory of the system enters this set, we assume that the system will operate safely, as is typically assumed in reach-avoid problems [39].
Remark 3.
It is worth noting that these expert controllers are typically designed with a primary focus on safety and may therefore be overly conservative in terms of performance optimization (i.e., minimizing the infinite-horizon cost (2)). By introducing the SACBF framework and the optimization-based controller (3), we can achieve near-optimal performance while preserving safety.
@@
One relevant question is whether the robust optimization problem (14) is actually feasible. In the following proposition, we address this concern by presenting a converse SACBF theorem. This result shows that if a safe policy exists (Assumption 1), then an SACBF must also exist. Consequently, the feasibility of problem (14) is guaranteed. Moreover, the proposition establishes that any optimal solution to (14) yields a valid SACBF with a non-empty safe set.
@@
Under Assumption 1, there exists a $\beta\in[0,1)$ such that the problem (14) is feasible and such that any optimal solution pair $(Q^{B*},q^{*})$ ensures that $Q^{B*}$ is an SACBF with a non-empty safe set $\mathcal{S}^{*}_{Q}=\{x\in\mathcal{X}|q^{*}(x)\leq 0\}$.
@@
$\mathcal{S}_{0}$ $=\left\{x\in\mathcal{X}|q(x)\leq 0\right\},$
$\partial\mathcal{S}_{0}$ $=\left\{x\in\mathcal{X}|q(x)=0\right\},$
$\operatorname{Int}(\mathcal{S}_{0})$ $=\left\{x\in\mathcal{X}|q(x)<0\right\}$ (15)
@@
With the above definitions, we follow the proof of the Converse CBF Theorem in the continuous time domain [2, Proposition 3] to prove the feasibility of (14). In particular, for any continuous $q$ such that (14b) and (4.2) hold, define a function $\alpha(\cdot):[0,\infty)\to\mathbb{R}$ by
@@
When $x$ is such that $q(x)=0$, i.e., $x\in\partial\mathcal{S}_{0}$, by the invariance of $\mathcal{S}_{0}$, we have $q(f(x,\pi_{\mathrm{s}}(x)))\leq 0$, and consequently $\alpha(0)\leq 0$. Meanwhile, by taking the extreme values 0 and r-r of $q(f(x,\pi_{\mathrm{s}}(x)))$ and $q(x)$ respectively, we have $\min_{r>0}\frac{\alpha(r)}{r}\leq\min_{r>0}\frac{0-(-r)}{r}=1$, which, combined with $\alpha(0)\leq 0$, implies that there exists a class $\mathcal{K}$ function $\bar{\alpha}$ upper-bounding $\alpha$ and satisfying
@@
with $x^{+}=f(x,\pi_{\mathrm{s}}(x))$. This implies that the new function $q^{\prime}$ satisfies all the constraints in (14). Moreover, since the zero sub-level set of $\max\{h,\;\bar{q}\}$ is non-empty, the zero sub-level set of $q^{\prime}$ is non-empty as well. Furthermore, it strictly holds that $\int_{x\in\mathcal{X}}q^{\prime}(x)dx<\int_{x\in\mathcal{X}}q^{*}(x)dx$ because (i) $q^{\prime}(x)\leq q^{*}(x),\;\forall x\in\mathcal{X}$, and (ii) $q^{*}>0$ and $q^{\prime}\leq 0$ in the zero sub-level set of $q^{\prime}$. These arguments prove that the new function $q^{\prime}$ provides a better solution than $q^{*}$, which contradicts the optimality of $q^{*}$.
@@
For the optimization problem (14), we replace the infinite sets $\mathcal{S}_{0}$, $\mathcal{X}$, and $\mathcal{S}_{0}\times U$ by the finite sets $\mathcal{S}_{\mathrm{s}}$, $\mathcal{X}_{\mathrm{s}}$, and $\mathcal{S}_{\mathrm{s}}\times U_{\mathrm{s}}$, respectively. Here $U_{\mathrm{s}}$ is the set of inputs sampled in $U$ through some sampling strategy [^2] We do not assume i.i.d. samples.. As a result, (14) is simplified to a constrained optimization problem with a finite number of constraints. If $Q^{B}_{\omega}$ and $q$ are linear in $\omega$, the problem (14) becomes linear and can be solved efficiently. In a more general case such as when these functions are represented by deep neural networks, a tractable solution involves transforming the constraints into soft penalties and adding them into the objective.
Note that the proposed method reduces conservatism through its optimization objective, because the objective of minimizing $q$ contributes to enlarging the safe set.
@@
The above sampling-based approach does not guarantee the validity of the approximated solution. [44], which studies learning CBFs, suggests leveraging the Lipschitz properties of $f$ and the approximated solution for principled constraint tightening. However, estimating the Lipschitz constants when $Q^{B}$ is parametrized by a neural network is non-trivial, and obtaining non-conservative Lipschitz constants for the unknown function $f$ is questionable. To address this limitation, we will propose a new tightening approach that exploits the inherent robustness of the SACBF. This will be elaborated on in Section V.
4.3 Supervised learning SACBFs from CBFs
At the end of this section, we present the most sample-efficient learning-based method for obtaining an SACBF, based on a rather restrictive assumption that we have a CBF at hand.
Assumption 2.
A CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is known.
In certain applications, such as adaptive cruise control [2], CBFs can be manually designed based on state constraints, such as the distance between adjacent vehicles. Moreover, there are mature techniques for synthesizing valid CBFs without model information [46, 45, 55]. However, as noted in the introduction, such CBFs cannot be directly used to design a safety filter when the model is not fully known.